OpenStudy (help_people):

Given the following functions f(x) and g(x), solve for (f ⋅ g)(2) and select the correct answer below: f(x) = 3x2 + 2 g(x) = x − 8 −42 −84 42 84

OpenStudy (help_people):

@Vocaloid @DaBest21

OpenStudy (anonymous):

im here

OpenStudy (anonymous):

first plug 2 in for x in g(x) so do 2-8

OpenStudy (anonymous):

What do u get @help_people

OpenStudy (help_people):

-6

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

in 3x2 is the 2 an exponent

OpenStudy (help_people):

yes

OpenStudy (anonymous):

ok in the future just do 3x^2

OpenStudy (help_people):

ok srry

OpenStudy (anonymous):

sokay

OpenStudy (anonymous):

now do 3*2^2+2

OpenStudy (anonymous):

what do u get

OpenStudy (anonymous):

remeber do 2^2 first

OpenStudy (help_people):

14

OpenStudy (anonymous):

yep so now do 14*-6 what do u get

OpenStudy (help_people):

-84

OpenStudy (anonymous):

yep thats ur answer

OpenStudy (anonymous):

Don't forget to medal and fan if you already haven't. Plus send me a message if you need more help @help_people im actually gonna take a break

OpenStudy (help_people):

wait could you help me w. one more please please please?

OpenStudy (help_people):

please ?

OpenStudy (help_people):

i will post it here so it does not take long

OpenStudy (anonymous):

k

OpenStudy (help_people):

Given the functions k(x) = 5x − 8 and p(x) = x − 4, solve k[p(x)] and select the correct answer below. k[p(x)] = 5x − 12 k[p(x)] = 5x − 28 k[p(x)] = 5x2 − 12 k[p(x)] = 5x2 − 28

OpenStudy (help_people):

the 2s in the last 2 answer choices are exponents

OpenStudy (anonymous):

well do (x-4)*5-8 5x-20-8 5x-12

OpenStudy (anonymous):

there u go

OpenStudy (anonymous):

i gotta take a break

OpenStudy (help_people):

thank you so much have fun you deserve it :)